OpenStudy (karatechopper):

5x+7y=-1 -2y+3z=9 7x-z=27 So I got a bit lost on this problem. Can someone walk me through?

OpenStudy (solomonzelman):

what methods do you know? Do you know how to use matrix substitution add or subtract the equations from each other

OpenStudy (karatechopper):

Uhm for this one its either substitution or elimination

OpenStudy (karatechopper):

No matrix

OpenStudy (austinl):

Ok, I have a thought... You can solve for z in the last equation, and then solve for y in the first equation and then you should be able to solve for x by substituting into the second equation.

OpenStudy (solomonzelman):

you can use matrix too, if you know how to. That's what I would suggest.

OpenStudy (karatechopper):

My teacher said not to do matrix plus I haven't learnt it yet from her. We learn that tomorrow. But I have to know how to do this cuz I have a quiz tomorrow

OpenStudy (solomonzelman):

Oh, ok.

OpenStudy (karatechopper):

Lemme tell yall how she told me to go about with the steps of solving.

OpenStudy (solomonzelman):

Then do substitution,

OpenStudy (solomonzelman):

5x+7y=-1 -2y+3z=9 7x-z=27 (recalling)

OpenStudy (karatechopper):

1)pick a variable to eliminate 2) pick a pair of equations and eliminate that variable 3)pick a different pair of equations and eliminate the same variable. 4)solve the two resulting equations

OpenStudy (solomonzelman):

Yup

OpenStudy (solomonzelman):

eliminate z in the last equation and substitute into the second one.

OpenStudy (karatechopper):

Do I not pick a pair twice tho?

OpenStudy (solomonzelman):

you will result with 2 variables after you substitute the z from the last equation to the second

OpenStudy (karatechopper):

Like, I did elimination and took out the y in the 1st and 2nd equation, then in the 2nd and 3rd equation I eliminated the y. Then I was confused where to go next

OpenStudy (karatechopper):

yes correct, I resulted in 2 variables

OpenStudy (solomonzelman):

tell me what exactly what you get, write it out.

OpenStudy (karatechopper):

Ok.

OpenStudy (solomonzelman):

I mean when you get 2 variables....

OpenStudy (austinl):

\(~~~~5x+7y=-1\\ -2y+3z=9\\ ~~~~~~7x-z=27\) We have this correct? Well we can take the third line, and solve for z. \(7x-z=27\) \(-z=-7x+27\) \(z=7x-27\) Then we can take the first line and solve for y. \(5x+7y=-1\) \(7y=-5x-1\) \(\displaystyle y=-\frac{5}{7}x-\frac{1}{7}\) Then we should be able to substitute it back into the second equation. \(-2y+3z=9\) \(\displaystyle -2(-\frac{5}{7}x-\frac{1}{7})+3(7x-27)=9\) Now we have all x's in this line and can therefore solve for x right? After you have that you should be able to easily solve out for the remaining variables.

OpenStudy (solomonzelman):

@austinL, why are you doing his work for him?

OpenStudy (austinl):

It was significant enough that I deemed it more appropriate to do it this way so that he can see it uninterrupted.

OpenStudy (karatechopper):

1st step 5x+7y=-1 -2y+3z=9 10x+14y=-2 + 21z-14y=63 ------------- 10x+21z=61 2nd step -2y+3z=9 7x-z=27 -2y+3z=9 + 21x-3z=81 ----------- -2y+21x=90

OpenStudy (karatechopper):

Those are the steps I have gone through so far..now I am confused where to take those two equations from there.

OpenStudy (solomonzelman):

let me check....

OpenStudy (karatechopper):

The next step is to solve the resulting equations...how to do that?!

OpenStudy (austinl):

The thing is, there are 3 variables but only 2 variables in each line making elimination method difficult to use.

OpenStudy (solomonzelman):

My bad it is -2

OpenStudy (karatechopper):

-2+63=61 lol

OpenStudy (solomonzelman):

Yeah, you are right!

OpenStudy (karatechopper):

So then how do I use substitution after my steps? @austinL

OpenStudy (karatechopper):

alright where do I take it from there then? @SolomonZelman

OpenStudy (solomonzelman):

I keep losing the equations b/c of the replies, Ill right them on my paper...

OpenStudy (karatechopper):

Alright

OpenStudy (austinl):

Go back and look at my steps, that is how I would set about solving this problem.

OpenStudy (karatechopper):

I see you do substitution before, Yeah I need to learn it my teachers way..

OpenStudy (solomonzelman):

look at your first equation (of the initial once, that are in your question)

OpenStudy (karatechopper):

5x+7y=-1

OpenStudy (solomonzelman):

ok. and the one that you got -2y+21x=90

OpenStudy (karatechopper):

ok.

OpenStudy (solomonzelman):

( I agree that substitution is easier though...)

OpenStudy (karatechopper):

Wait, what about me other equation!!!

OpenStudy (austinl):

But I fail to see how this(your teachers method) can work with this problem. It works just dandy for systems set up like. ax+by+cz=d ex+fy+gz=h ix+jy+kz=l yours is set up, ax+by=c -dy+ez=f gx-hz=i

OpenStudy (karatechopper):

Right, I agree. Somehow she made it work tho..

OpenStudy (austinl):

You said this is for a quiz tomorrow right? Does the REQUIRE you to use the method you mentioned above?

OpenStudy (solomonzelman):

it probably would, I like substitution and matrix better....

OpenStudy (solomonzelman):

karatechopper, it would work better with 4 equations, but 3 equations like these ones, Nooooooooooo

OpenStudy (karatechopper):

Yes. it requires it

OpenStudy (solomonzelman):

Normally you subtract one equation from both other equations to get rid of a variable completely...

OpenStudy (austinl):

\( 5x+7y=-1\) \(-2y+3z=9\) \(7x-z=27\) This is what we have. Lets take the first two lines and eliminate y, \(5x+7y=-1\) \(-2y+3z=9\) \(2(5x+7y=-1)\) \(7(-2y+3z=9)\) This will give, \(10x+14y=-2\) \(\underline{-14y+21z=63}\) \(10x+21z=61\) Then we can take that and the last line, \(10x+21z=61\) \(7x-z=27\) You can then eliminate one variable and solve for one of them. You can then substitute back in what you have easily to solve for the remaining variables.

OpenStudy (austinl):

Anywho, I have to go now, take care guys!

OpenStudy (karatechopper):

oh..i see..

OpenStudy (solomonzelman):

@austinL is right! Can you do it from there?

OpenStudy (solomonzelman):

yes you could.......

OpenStudy (karatechopper):

I believe so yes!

OpenStudy (solomonzelman):

are you ready for a quiz?

OpenStudy (karatechopper):

Not really. I need for practice on these problems.

OpenStudy (karatechopper):

Would you give me a quiz? I need to practice problems like this one. I think I am good with the 3 variables in each equation one